OpenStudy (anonymous):

what is f^-1(4) when f(x)=(3x-4)/5?

OpenStudy (anonymous):

@ranga help?

OpenStudy (anonymous):

@ParthKohli help?

OpenStudy (ranga):

You need to find the inverse of f(x). To do that follow these steps: 1. Replace f(x) with y. 2. Interchange x and y. That is, put x in the place of y and y in the place of x. 3. Solve for y. This will be the inverse. 4. replace y with f^-1. 5. Find f^-1(4) by putting x = 4 in the inverse function.

OpenStudy (anonymous):

so f^-1(4) would = -4?

OpenStudy (ranga):

I m getting a different answer. If you show me the steps I can try to spot any mistakes in the steps.

OpenStudy (anonymous):

i dont know how much is f^-1(4)

OpenStudy (ranga):

First you have to find the inverse of the function. I gave you the steps above. What is the inverse function you are getting? f^-1(x) = ?

OpenStudy (anonymous):

3y-4/5?

OpenStudy (ranga):

What you have is step #2. x = (3y - 4) / 5. In step#3, solve for y. That is y = ?

OpenStudy (anonymous):

how would i solve for y ?

OpenStudy (ranga):

x = (3y - 4) / 5 To solve for y, isolate y. multiply both sides by 5. Add 4 to both sides. divide both sides by 3.

OpenStudy (anonymous):

i cant do it i nee to see step by step :/

OpenStudy (ranga):

x = (3y - 4) / 5 multiply both sides by 5. What do you get?

OpenStudy (anonymous):

so it would be x=15y-20?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

When you multiply the left hand side by 5, x will become 5x. When you multiply the right hand side by 5, it will cancel out the 5 in the denominator and you will be left with just (3y - 4)

OpenStudy (anonymous):

ooh alright got it now whats next? replace y with f^-1 so it'll be 5x=3f^-1-4?

OpenStudy (ranga):

Not yet. You have to solve for y first. x = (3y - 4) / 5 multiply both sides by 5: 5x = 3y - 4 add 4 to both sides: 5x + 4 = 3y divide both sides by 3 (5x+4)/3 = y or y = (5x + 4) / 3 Now do step 4 and replace y with f^-1(x) f^-1(x) = (5x + 4) / 3 f^-1(4) = ?

OpenStudy (anonymous):

-4 @ranga

OpenStudy (anonymous):

Then just plug in -4 into the x?

OpenStudy (ranga):

Not -4, plug in +4.

OpenStudy (ranga):

What do you get when you put x = 4 in (5x + 4) / 3 ?

OpenStudy (anonymous):

i get 8

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

so thats the answer?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

thank you @ranga you're the best!

OpenStudy (ranga):

You are welcome. Thanks.